THE OPTICAL STARK SPECTRUM OF THE [11.9]Ω = 3/2 − X^{3}Π_{3/2} BAND SYSTEM OF PLATINUM MONOFLUORIDE, PtF

\underline{CHENGBING QIN} AND TIMOTHY C. STEIMLE , Department of Chemistry and Biochemistry, Arizona State University, Tempe,AZ 85287.

Recently the O’Brien group^{a} reported on the field-free detection and analysis of the [11.9]= 3/2 − X^{2}Π_{3/2} band system of PtF using intracavity laser absorption. The hollow cathode condition limited the spectral resolution, which was insufficient to fully resolved the hyperfine splitting of ^{1}^{9}^{5}Pt (I=1/2) and ^{1}^{9}F (I=1/2). Here we report laser-induced fluorescence spectra of the 1-0 and 0-0 bands of the [11.9]= 3/2 − X^{2}Π_{3/2} transition of PtF, obtained at near natural linewidth resolution (FWHM 40MHz) using supersonic molecular beam techniques. The spectra show a complex, cleraly resolved hyperfine structure which has siginificant contributions from magnetic term in ^{1}^{9}^{5}Pt and ^{1}^{9}F. The spectra of 1^{9}^{4}PtF, ^{1}^{9}^{5}PtF, ^{1}^{9}^{6}PtF, ^{1}^{9}^{8}PtF isotopogues have been assigned and analyzed. The electric field induced dependence of the R(1.5) branch of the ^{1}^{9}^{5}PtF isotopogue was analyzed to produce permanent electric dipole moment, μ, of 3.45 D and 2.32 D for the X^{2}Π_{3/2} and [11.9]Ω = 3/2 states, respectively. A comparision with Ab Initio prediction^{b}^{,}^{c} will be given.

aK.G. Handler, R.A. Harris, L.C. O’Brien and J.J.O’Brien J.Mol. Spectrosc. 265 39, 2011.

bW. Liu and R. Franke J. Comput. Chem. 23 564, 2001.

cW. Zou, Y. Liu and J. E. Boggs Dalton Transactions, 39 2023, 2000